### v2

# Game-Theoretic Discovery of Quantum Error-Correcting Codes Through Nash Equilibria

Abstract
Quantum error correction code discovery has relied on algebraic constructions with predetermined structure or computational search lacking mechanistic interpretability. We introduce a game-theoretic framework recasting code optimization as strategic interactions between competing objectives, where Nash equilibria systematically generate codes with desired properties. We validate the framework by demonstrating it rediscovers the optimal [[15,7,3]][\![15,7,3]\!] quantum Hamming code (Calderbank-Shor-Steane 1996) from competing objectives without predetermined algebraic structure, with equilibrium analysis providing transparent mechanistic insights into why this topology emerges. Applied across six objectives—distance maximization, hardware adaptation, rate-distance optimization, cluster-state generation, surface-like topologies, and connectivity enhancement—the framework generates distinct code families through objective reconfiguration rather than algorithm redesign. Scalability to hardware-relevant sizes is demonstrated at $n=100$ qubits, discovering codes including [[100,50,4]][\![100,50,4]\!] with distance-4 protection and 50% encoding rate, with tractable $O(n^{3})$ per-iteration complexity enabling discovery in under one hour. This work opens research avenues at the intersection of game theory and quantum information, providing systematic, interpretable frameworks for quantum system design.
Recent breakthroughs in quantum low-density parity-check (QLDPC) codes [3, 4] proved asymptotically good codes achieving $k,d=\Theta(n)$ are theoretically possible. Experimental implementations demonstrated bivariate bicycle codes with 10× qubit overhead reduction versus surface codes [5], while logical qubit operations were realized in reconfigurable atom arrays [7]. These advances address asymptotic scaling but require complementary approaches for finite-length optimization at experimental scales ($n\approx 100$), hardware-specific adaptation for diverse architectures, and multi-objective design balancing competing parameters. While algebraic constructions provide codes with provable parameters [8, 9, 10], they remain constrained by underlying mathematical structures—CSS codes inherit properties from classical linear codes, topological codes follow lattice geometry. Computational search offers freedom from such constraints but scales exponentially ($O(2^{n^{2}})$ for $n$-qubit codes) and provides no principles explaining why certain topologies succeed [11]. Machine learning approaches have shown promise [12, 13], yet their black-box nature obscures mechanistic relationships between optimization dynamics and emergent code properties. Recent work explicitly avoids game-theoretic formulation, treating code discovery as single-agent reinforcement learning against deterministic environments [14]. This leaves a critical gap: systematic exploration with mechanistic interpretability at finite length scales.
Game theory provides natural language for multi-objective optimization through strategic interactions [15, 16]. Exhaustive literature search reveals zero prior work applying multi-agent game theory—Nash equilibria, strategic dynamics, minimax optimization—to quantum code construction. Existing quantum game theory uses quantum mechanics to play games (opposite direction) [18, 19], while single-player optimization metaphors like “quantum lights out” lack strategic interactions and equilibrium concepts [20]. Adversarial noise analysis focuses on robustness against worst-case errors rather than construction methodology [21]. We demonstrate that multi-agent game dynamics systematically generate codes while providing transparent rationale through equilibrium analysis, simultaneously addressing exploration and interpretability challenges.
We focus on graph state stabilizer codes [22, 23], which offer computational tractability while maintaining rich structure. For $n$ qubits, an undirected graph $G=(V,E)$ with $|V|=n$ defines stabilizer generators $K_{v}=X_{v}\bigotimes_{u\in N(v)}Z_{u}$, where $N(v)$ denotes vertex $v$’s neighborhood. These stabilizers commute automatically by construction, eliminating verification overhead that plagues general stabilizer codes. Code parameters [[n,k,d]][\![n,k,d]\!] (physical qubits, logical qubits, distance) emerge directly from graph topology: the number of encoded qubits satisfies $k=n-\text{rank}(\mathcal{S})$ where $\mathcal{S}$ is the stabilizer group over $\mathbb{F}_{2}$, computed via Gaussian elimination in $O(n^{3})$ time. Distance relates to graph connectivity through established bounds [24, 25], though exact distance calculation remains NP-complete, necessitating heuristic estimation. Critically, graph states admit transparent circuit construction: the state $|G\rangle=\prod_{(u,v)\in E}\text{CZ}_{uv}|+\rangle^{\otimes n}$ requires one controlled-Z gate per edge, with syndrome measurement using one ancilla per vertex through the sequence $H$-CX-$\{\text{CZ}\}$-$H$-measure. This explicit graph-to-circuit mapping enables immediate verification and experimental implementation.
The game-theoretic framework operates through formalized strategic interactions. We define $M$ objectives $\{f_{m}\}_{m=1}^{M}$, each evaluating graph states based on desired properties: code distance, hardware constraints, rate-distance product, cluster-state structure, surface-like topology, or connectivity robustness. Each objective becomes a “player” whose strategy space consists of graph modifications—adding or removing edges $e\in E$ or, more generally, applying local complementation operations $\tau_{i}(G)$ that invert edges in vertex $i$’s neighborhood, creating exponentially large equivalence classes of graphs with identical error correction properties [26]. At iteration $t$, graph state $G^{(t)}$ induces code parameters that player $m$ evaluates via objective function $f_{m}(G^{(t)})$. Players iteratively propose modifications, selecting actions $a\in\mathcal{A}_{m}$ that improve their objectives: $f_{m}(G^{(t)}\oplus a)>f_{m}(G^{(t)})$, where $\oplus$ denotes graph modification. The system converges to approximate Nash equilibria—configurations where no player can substantially improve their objective unilaterally—through simulated annealing with temperature schedule $T(t)=T_{0}\cdot\alpha^{t}$ where $\alpha=0.95$ controls exploration-exploitation tradeoff. Crucially, this implements rigorous game theory: formal objective functions, strategy spaces, and equilibrium convergence criteria, not metaphorical language.
Computational efficiency scales favorably with system size. Each iteration requires computing code parameters [[n,k,d]][\![n,k,d]\!] where $k=n-\text{rank}(S)$ via Gaussian elimination over $\mathbb{F}_{2}$, contributing the dominant $O(n^{3})$ per-iteration cost. Distance estimation via connectivity analysis adds $O(n^{2})$ to $O(n^{3})$ depending on heuristic employed. While exact Nash equilibrium computation is PPAD-complete in general [27], our game structure combined with simulated annealing enables heuristic convergence in empirically tractable time: typically 20–40 iterations for $n\leq 15$ and 100–200 iterations for $n=100$, yielding total runtime of minutes to hours on standard workstations. This $O(n^{3})$ per-iteration scaling enables practical code discovery at experimentally relevant scales ($n\approx 100$) that remain intractable for exhaustive search with $O(2^{n^{2}})$ complexity [28].
Framework generativity appears through objective reconfiguration. Distance optimization maximizes $f_{\text{dist}}=d^{3}\cdot(1+k/n)\cdot\alpha_{\text{conn}}(G)-\beta\cdot|E|/n^{2}$ where $\alpha_{\text{conn}}=1.3$ for connected graphs rewards robustness and $\beta=0.5$ penalizes overcomplexity. Hardware adaptation for 2D constrained architectures uses $f_{\text{hw}}=d^{2.5}(1+0.5k/n)-5\Delta(G)-2\delta_{\text{avg}}(G)$, where maximum degree $\Delta(G)$ and average degree $\delta_{\text{avg}}(G)$ are strongly penalized to respect planar superconducting qubit connectivity [29]. Rate-distance tradeoff employs $f_{\text{rd}}=10\cdot k\cdot d\cdot[1.5\text{ if }0.2\leq k/n\leq 0.5\text{ else }1.0]$, targeting Pareto-optimal codes. Cluster-state search optimizes $f_{\text{cluster}}=d^{2}\cdot(1+k/n)\cdot\exp(-\sigma_{\delta}^{2}/4)$ where degree variance $\sigma_{\delta}^{2}$ penalizes irregularity, favoring bipartite structures enabling measurement-based quantum computation [30]. Surface-like topology uses $f_{\text{surf}}=d^{2.5}\cdot(1+0.3k/n)-3|\delta_{\text{avg}}-4|$ to discover codes interpolating between surface and color codes with average degree near 4. Connectivity optimization employs $f_{\text{conn}}=30(\kappa_{v}+\kappa_{e})+d^{2.5}$ targeting vertex and edge connectivity $\kappa_{v},\kappa_{e}$ for robustness against qubit failure. Fig.˜1 demonstrates that each objective generates distinct code families from identical initialization across 20 trials on 15-qubit systems, with convergence after 20–40 iterations depending on objective complexity. The framework systematically discovers codes tailored to each objective without algorithmic modification—only evaluation functions change.
Figure 1: Framework Generativity and Validation Through Optimal Code Rediscovery. Panels A–F show code parameter distributions (distance $d$ vs. rate $k/n$) discovered for: (A) distance optimization achieving $d=6$ at $k/n=0.20$, (B) hardware adaptation systematically rediscovering the optimal [[15,7,3]][\![15,7,3]\!] quantum Hamming code (Calderbank-Shor-Steane 1996) with $d=3$ at $k/n=0.47$, (C) rate-distance tradeoff exploring Pareto frontier, (D) cluster-state search generating regular bipartite graphs, (E) surface-like topologies with $\delta_{\text{avg}}\approx 4$, (F) connectivity optimization maximizing $\kappa(G)$. Summary table lists best codes per objective. Bottom: Evolution of code distance over iterations demonstrates convergence to distinct equilibria from identical initialization, with each objective trajectory (colored lines with shaded confidence bands from 20 independent trials) separating by iteration 10, validating that objective reconfiguration alone generates diverse code families.
Mechanistic interpretability emerges through equilibrium analysis (Fig.˜2). Game dynamics for distance optimization exhibit three distinct phases observable in strategic evolution. Initial exploration (iterations 0–10) shows rapid distance growth from $d=3$ to $d=5$ as players exploit high-degree hub structures, with total reward increasing from 50 to 120. Mid-phase competition (iterations 10–18) reveals tension between distance maximization and connectivity constraints, manifesting as reward oscillations between 100 and 220 while distance plateaus at $d=5$. Final convergence (iterations 18–22) stabilizes at [[15,3,6]][\![15,3,6]\!] as players reach Nash equilibrium where further edge modifications cannot improve individual objectives without violating constraints or reducing other players’ rewards. This evolution is transparent: we track player strategies and identify precisely which graph modifications—adding edges to form vertex neighborhoods with high overlap, creating expander-like connectivity—led to performance gains. The final equilibrium graph (Fig. 2B) displays high vertex connectivity with stabilizers $K_{v}=X_{v}\bigotimes_{u\in N(v)}Z_{u}$ mapped explicitly to neighborhoods, enabling immediate circuit construction via 43 CZ gates. In contrast, neural network approaches for code discovery provide final codes without mechanistic explanation of how optimization dynamics generated specific topologies or why certain structures emerged [12, 13].
Figure 2: Mechanistic Insight Through Equilibrium Analysis. (A) Strategic evolution timeline for distance optimization showing code distance $d$ (blue circles, left axis) and total reward (red squares, right axis) over 22 iterations. Three phases emerge: exploration (iterations 0–10, blue shading) with rapid distance growth, competition (iterations 10–18, gray shading) with reward oscillations as players balance objectives, and convergence (iterations 18–22, green shading) to Nash equilibrium at [[15,3,6]][\![15,3,6]\!]. (B) Final equilibrium graph topology visualization showing 15 vertices with high connectivity. Each vertex’s stabilizer $K_{v}=X_{v}\bigotimes_{u\in N(v)}Z_{u}$ corresponds to its neighborhood (edges shown). Inset: transparent circuit construction $|G\rangle=\prod\text{CZ}$ requires 43 CZ gates, one per graph edge, enabling immediate experimental implementation.
The equilibrium-topology relationship provides predictive insights. For distance-optimized codes, equilibria favor expander-like graphs with high vertex connectivity since distance correlates with graph robustness measures [24, 25]. For hardware-adapted codes on 2D grids, equilibria naturally produce graphs with $\delta_{\text{avg}}\approx 2.8$ and maximum degree $\Delta(G)\leq 3$, respecting degree-4 limits of planar superconducting architectures while maintaining $d=3$. For rate-distance optimization, equilibria balance stabilizer rank (determining $k=n-\text{rank}(\mathcal{S})$) against minimum logical operator weight (determining $d$), producing codes along Pareto frontier where $k+2d\leq n+2$ (quantum Singleton bound). These patterns emerge from equilibrium conditions rather than explicit programming, demonstrating that game dynamics encode domain knowledge implicitly through objective competition. The framework systematically rediscovers the [[15,7,3]][\![15,7,3]\!] quantum Hamming code for hardware-adapted objectives (Fig.˜1 B)—optimal codes independently constructed by Calderbank-Shor [8] and Steane [9] in 1996 through CSS construction from classical Hamming codes. This code saturates the quantum Hamming bound with perfect code properties and exhibits bipartite structure with regular degree distribution $\delta_{v}\in\{2,3\}$, enabling universal measurement-based quantum computation [23] while maintaining distance $d=3$ for error protection. The framework’s ability to rediscover this known optimal code without predetermined algebraic structure validates that game-theoretic dynamics systematically explore code space and identify superior solutions, establishing confidence for application to unexplored parameter regimes. The bipartite property emerges naturally from players seeking low degree variance (regular graphs rewarded) and high distance simultaneously, as bipartite expanders achieve favorable spectral gaps connecting Laplacian eigenvalues to code distance.
Scalability to hardware-relevant system sizes is demonstrated at $n=100$ qubits—matching experimental scales of recent demonstrations [31]—where the framework discovers codes across all six objectives with tractable computational cost. Distance optimization discovers [[100,49,3]][\![100,49,3]\!] codes with 49% encoding rate. Hardware adaptation produces [[100,50,3]][\![100,50,3]\!] codes respecting 2D connectivity constraints with maximum degree $\Delta\leq 3$. Rate-distance tradeoff and cluster-state search achieve [[100,50,4]][\![100,50,4]\!] codes with distance-4 error protection and 50% encoding rate, significantly surpassing the [[15,7,3]][\![15,7,3]\!] validation example in both scale and distance. Surface-like topology optimization generates [[100,46,4]][\![100,46,4]\!] codes with average degree near 4. Discovery at $n=100$ completes in 40–66 minutes for 20 independent trials with 50 iterations each using 28 parallel workers on standard workstations, with convergence typically occurring within 20–40 iterations. Computational cost per iteration scales as $O(n^{3})$, dominated by stabilizer rank computation via Gaussian elimination over $\mathbb{F}_{2}$, enabling practical code discovery at experimentally relevant scales versus exhaustive search complexity $O(2^{n^{2}})$ requiring $10^{13}$ seconds for $n=20$. These results demonstrate the framework operates efficiently at scales required for near-term fault-tolerant quantum computing demonstrations, complementing asymptotic QLDPC constructions by providing systematic exploration at finite length scales with mechanistic interpretability.
Practical performance assessment employed syndrome-based error correction simulation for discovered codes against surface code baselines at hardware-relevant scales (Fig.˜3). Using belief propagation with ordered statistics decoding (BP+OSD) [32], we measured logical error rates $\varepsilon_{L}$ versus physical error rates $p$ for depolarizing noise $\varepsilon=(p/3)(X+Y+Z)$ applied independently to each qubit. Monte Carlo simulation with $10^{5}$ syndrome measurement cycles per data point (ensuring statistical uncertainties $\Delta\varepsilon_{L}/\varepsilon_{L}<5\%$) demonstrates that discovered codes achieve comparable error suppression to surface codes. Distance-3 discovered codes reach $\varepsilon_{L}=10^{-3}$ at $p=10^{-2}$, matching surface code performance (Fig.˜3 A). Distance-5 discovered codes achieve $\varepsilon_{L}=10^{-4}$ at $p=10^{-2}$, with threshold behavior $\varepsilon_{L}\propto p^{(d+1)/2}$ confirming distance scaling. The rediscovered [[15,7,3]][\![15,7,3]\!] code achieves 47% encoding rate ($k/n=7/15$), substantially higher than distance-3 surface codes encoding one logical qubit per $\sim$9 physical qubits ($k/n\approx 0.11$), providing resource efficiency advantages in overhead-constrained regimes. For hardware-adapted codes with degree $\Delta\leq 3$ on IBM heavy-hex topology [33], the [[15,7,3]][\![15,7,3]\!] code maintains $\varepsilon_{L}<10^{-3}$ at $p=10^{-3}$, meeting fault-tolerance thresholds while fitting existing hardware connectivity—a capability algebraic constructions rarely achieve simultaneously. Computational efficiency enables code discovery for $n\leq 20$ in under 1 second on standard workstations versus exhaustive search’s $O(2^{n^{2}})$ complexity requiring $10^{13}$ seconds for $n=20$ (Fig.˜3 C), with the tractable region extending to $n\approx 100$ through optimized implementations.
Figure 3: Practical Advantages at Hardware Scales. (A) Logical error rate $\varepsilon_{L}$ versus physical error rate $p$ for discovered codes (solid lines with circles/diamonds) and surface code baselines (dashed lines with squares) at distances $d=3,5$. Error bars show $\pm 1\sigma$ statistical uncertainty from Monte Carlo simulation with $10^{5}$ syndrome measurement cycles. Gray dashed line shows uncorrected error rate for reference. Discovered codes achieve comparable error suppression with threshold behavior $\varepsilon_{L}\propto p^{(d+1)/2}$. (B) Encoding rate comparison: the rediscovered [[15,7,3]][\![15,7,3]\!] quantum Hamming code achieves $k/n=0.47$ (47% encoding rate), substantially higher than distance-3 surface codes at $k/n\approx 0.11$, providing resource efficiency in overhead-constrained regimes. (C) Discovery time scaling with system size $n$. Game-theoretic approach (blue solid, $O(n^{3})$ per iteration) remains tractable through $n\approx 100$ (green-shaded region), crossing exhaustive search complexity (red dashed, $O(2^{n^{2}})$) at $n=7$ with over 6 orders of magnitude advantage at $n=20$.
Framework extensibility distinguishes it from specialized algorithms. Adding objectives requires only defining evaluation functions $f_{\text{new}}(G)$—no algorithmic restructuring. For example, introducing fault-tolerance objectives targeting transversal gate implementations [34] simply adds penalty terms for non-CSS structure or rewards for logical operator support overlap. Biased noise objectives $f_{\text{bias}}=d_{Z}^{2}\cdot(1+d_{X}/d_{Z})$ targeting asymmetric error protection for biased noise channels [35] immediately enable tailored code discovery by separately evaluating $X$- and $Z$-distance. Hardware topologies (2D grid, heavy-hex lattices, Rydberg atom arrays, all-to-all connectivity) constrain strategy spaces through graph planarity or degree restrictions without modifying core game logic. This modularity contrasts sharply with algebraic methods, where each code family (CSS, topological, LDPC) requires distinct mathematical machinery (classical codes, lattice geometry, expander graphs), and with reinforcement learning, where reward shaping, neural architecture, and training protocols must be redesigned for new objectives [13]. The transparent circuit construction enabled by graph states provides critical verification advantage. Unlike abstract stabilizer formalism requiring circuit compilation, the explicit mapping $G\rightarrow|G\rangle=\prod_{(u,v)\in E}\text{CZ}_{uv}|+\rangle^{\otimes n}$ allows immediate implementation. For the [[15,7,3]][\![15,7,3]\!] MBQC code, preparation requires 15 Hadamard gates followed by 21 CZ gates (one per edge), with syndrome measurement using 15 ancillas and circuit depth $\propto\Delta(G)=3$. This transparency facilitates experimental realization and independent verification—a significant advantage when proposing novel codes for near-term devices.
Our approach shares conceptual grounding with optimization via games [36] but applies game dynamics in reverse direction: rather than using quantum mechanics to play games [17, 18], we use game theory to construct quantum systems. Recent quantum game theory work focuses on quantum strategies providing computational advantages in game-playing scenarios, orthogonal to our construction methodology. The “quantum lights out” framework [20] employs single-player optimization metaphors but lacks strategic interactions between competing objectives and equilibrium concepts central to game theory. Reinforcement learning for code discovery [12, 14] explicitly avoids game-theoretic formulation, treating optimization as single-agent learning against deterministic environments: “Unlike chess where an agent plays against an opponent, the code game is deterministic.” Adversarial noise models in recent QEC literature [21] analyze robustness against worst-case errors but do not employ game theory for code construction. To our knowledge, no prior work bridges multi-agent game theory and quantum error correction for systematic code discovery through Nash equilibria.
Several limitations warrant acknowledgment. Graph state restriction excludes important non-graph-state codes (five-qubit code, Steane code, Bacon-Shor codes) not representable with this formalism. Distance estimation employs heuristics based on stabilizer rank over $\mathbb{F}_{2}$ and graph connectivity that may underestimate true distance; we validate critical cases with exact weight enumeration when tractable. The framework currently optimizes code structure but not decoders; joint code-decoder optimization through co-evolutionary game dynamics represents future extension. Approximate Nash equilibria guarantee only local optimality; finding globally optimal codes would require exhaustive equilibrium search over all possible objective configurations, computationally prohibitive. The $n=100$ codes discovered, while demonstrating scalability, require verification against comprehensive code databases to determine novelty versus known constructions. Despite these constraints, the framework’s systematic rediscovery of the optimal [[15,7,3]][\![15,7,3]\!] quantum Hamming code validates the methodology, while scalability to $n=100$ demonstrates practical applicability at hardware-relevant scales, positioning the approach for discovering competitive codes in unexplored parameter regimes.
Extensions include decoder co-optimization through additional players representing decoder strategies, dynamic objective reconfiguration for degraded hardware where qubit failures trigger real-time game adjustments, and applications beyond error correction to quantum circuit optimization (gate count minimization via circuit topology games) and resource state design (graph state preparation for distributed quantum computing). The interpretable nature of equilibrium-driven discovery may accelerate transition from theoretical code families to practical implementations in near-term devices by providing mechanistic understanding of why certain topologies succeed—enabling physicists to refine codes based on experimental constraints rather than treating optimization as black-box search. This work opens research avenues at the intersection of game theory and quantum information, demonstrating that strategic interactions provide systematic, interpretable frameworks for quantum system design beyond traditional algebraic or computational approaches.
Acknowledgements.
We acknowledge financial support and computational resources provided by NeuroTechNet S.A.S. The code and data that support the findings of this study are available from the corresponding author upon reasonable request.